OpenStudy (anonymous):

Medalseason which angles are coterminal with 2pi/3 -13pi/3 -10pi/3 -4pi/3 5pi/3

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

i am trying to work it out ..

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

tbh i really cant work this out like the math i am coming with is not right super sorry for wasting your time :(

OpenStudy (thomas5267):

I am not sure but I think \( \dfrac{2\pi}{3}+2\pi n\) for all positive and negative integer n will be coterminal to \(\dfrac{2\pi}{3}\).

OpenStudy (anonymous):

oh ok thanks guys !